OpenStudy (anonymous):

is (1,1) a solution to 3y=2x+1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3(1) = 2(1)+1 3=3

Parth (parthkohli):

You can try substituting x with 1 and y with 1. \( \color{Black}{\Rightarrow 3(1) = 2(1) + 1}\) \( \color{Black}{\Rightarrow 3 = 2 + 1}\) \( \color{Black}{\Rightarrow 3 = 3 \checkmark }\)

OpenStudy (anonymous):

thanks